OpenStudy (anonymous):

collect the like terms and give full answer a squared + 3b squared - a squared + 2b squared collect the like terms and give full answer a squared + 3b squared - a squared + 2b squared @Mathematics

OpenStudy (asnaseer):

\[a^2+3b^2-a^2+2b^2\]so what you need to do is collect all terms involving a squared and, separately, all terms involving b squared. the terms involving a squared are:\[a^2-a^2=0\]and terms involving b squared are:\[3b^2+2b^2=5b^2\]putting it all together gives:\[a^2+3b^2-a^2+2b^2=5b^2\]